### v3

@@ 4.3 Construction in the case of general ${\bf D}$
${\bf G}_{1}=\left[\begin{array}[]{rr|r}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}&0\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}&-\Gamma_{1}\\ \hline\cr 0&\Gamma_{1}^{\top}&\Phi_{1}\end{array}\right],\qquad\Phi_{1}=-\Phi_{1}^{\top}.$
If $\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right]$ satisfies condition (66a), then in an analogous way we construct
$\widetilde{\bf T}_{1}=\mathop{\rm diag}\nolimits({\bf T}_{1},{\bf I},{\bf I}),\quad\widetilde{\bf V}_{1}=\mathop{\rm diag}\nolimits({\bf I},{\bf V}_{1},{\bf I})$
@@
${\bf G}_{2}=\widetilde{\bf T}_{1}\widetilde{\bf V}_{1}^{\top}{\bf G}_{1}\widetilde{\bf V}_{1}\widetilde{\bf T}_{1}^{-1}=\left[\begin{array}[]{rr|rr}-\widetilde{\bf A}_{2}&-\widetilde{\bf B}_{2}&0&0\\ \widetilde{\bf C}_{2}&\widetilde{\bf D}_{2}&-\Gamma_{2}&-\Gamma_{11}\\ \hline\cr 0&\Gamma_{2}^{\top}&\Phi_{2}&-\Gamma_{12}\\ 0&\Gamma_{11}^{\top}&\Gamma_{12}^{\top}&\Phi_{1}\end{array}\right]$
and we continue. Since the size of the matrices ${\bf T}_{j}$ is monotonically decreasing, this procedure will terminate after a finite number of $k$ steps with
@@
$\widetilde{\bf T}_{j}=\mathop{\rm diag}\nolimits({\bf T}_{j},{\bf I},{\bf I}),\quad\widetilde{\bf V}_{j}=\mathop{\rm diag}\nolimits({\bf I}_{\ell_{j}},{\bf V}_{j},{\bf I}),$
and $\ell_{j}$ is the size of ${\bf T}_{j}$ for $j=0,\ldots,k-1$.
@@
Remark 26.
The recursive procedure described above requires at each step the computation of three singular value decompositions in order to check the ranks of the matrices $\widetilde{\bf B}_{j}$ and $\widetilde{\bf C}_{j}$ and in order to construct bi-orthogonal bases so that (80) holds. While each step of this procedure can be implemented in a numerically stable way, the consecutive rank decisions make the aggregate procedure difficult to analyze, similar to the case of staircase algorithms ByeMX07 ; DemK93a ; DemK93b . In general the strategy should be adapted toward the goal of obtaining a realization in port-Hamiltonian form that is robust to small perturbations, see BanMNV20 ; MehV20 for some ways to do this.
@@ 4.4 Explicit solution of linear matrix inequalities via even pencils
We have seen that to check the existence of the transformation to port-Hamiltonian form and to explicitly construct the transformation matrices ${\bf T},{\bf V}$ in (8) is equivalent to considering the solution of the linear matrix inequality (9). In this subsection we combine the recursive procedure presented in the previous subsection with the construction of a staircase like form (Van79 ) for the even pencil (60).
@@
$\Psi_{1}({\bf Q}_{1})$ ::=:= $\left[\begin{array}[]{c|cc}-{\bf A}_{11}^{\top}{\bf Q}_{1}-{\bf Q}_{1}{\bf A}_{11}&-{\bf Q}_{1}{\bf A}_{12}-{\bf A}_{21}^{\top}{\bf Q}_{22}&{\bf C}_{31}^{\top}-{\bf Q}_{1}{\bf B}_{13}\\ \hline\cr-{\bf A}_{12}^{\top}{\bf Q}_{1}-{\bf Q}_{22}{\bf A}_{21}&-{\bf A}_{22}^{\top}{\bf Q}_{22}-{\bf Q}_{22}{\bf A}_{22}&{\bf C}_{32}^{\top}-{\bf Q}_{22}{\bf B}_{23}\\ {\bf C}_{31}-{\bf B}_{13}^{\top}{\bf Q}_{1}&{\bf C}_{32}-{\bf B}_{23}^{\top}{\bf Q}_{22}&{\bf S}_{1}\end{array}\right]$
=:absent:=:= : $\left[\begin{array}[]{cc}-{\bf A}_{1}^{\top}{\bf Q}_{1}-{\bf Q}_{1}{\bf A}_{1}&{\bf C}_{1}^{\top}-{\bf Q}_{1}{\bf B}_{1}\\ {\bf C}_{1}-{\bf B}_{1}^{\top}{\bf Q}_{1}&{\bf D}_{1}+{\bf D}_{1}^{\top}\end{array}\right]\geq 0$
@@
Note the above process is actually a special staircase form reduction process that deflates the singular part and higher index of the eigenvalue infinity of the even pencil (60), BruM08 ; ByeMX07 .
@@
If we consider the positive definite solution of the KYP equation, we can directly check if (106) has a deflating subspace associated with a set of $\ell$ finite eigenvalues chosen such that the deflating subspace is as in (60). If such a deflating subspace exists and ${\bf W}_{1}$ is invertible, then we can compute a Hermitian positive definite matrix ${\bf Q}_{k}$ associated with (106). We comment that an analogous characterization of solutions of singular H control problems via matrix pencils was given in CopS92 .
Remark 27.
Note that to check the passivity of (5) it is only necessary to have a positive semidefinite solution to (9). Thus, if one only wants to check passivity, then Part a) in Lemma 23 can be relaxed to $\mathsf{Ker}\,{\bf B}_{1}\subseteq\mathsf{Ker}\,{\bf C}_{1}^{\top}$. In this case the transformation to the form (97) can still be made, but ${\bf Q}_{22}$ in (98) is only positive semidefinite, and $\mathsf{Ker}\,{\bf Q}_{22}\subseteq\mathsf{Ker}\,{\bf Q}_{12}$. Then ${\bf Q}_{0}$ can still be defined but instead of ${\bf Q}_{11}^{-1}$ one needs to use the Moore-Penrose pseudoinverse, see GolV96 , of ${\bf Q}_{11}$. However, in this case ${\bf Q}_{0}$ and the resulting solution ${\bf Q}$ cannot be positive definite. Thus, in this situation, (5) may be a passive system that cannot be transformed to a standard port-Hamiltonian system of the form (6) system.
@@
Example 28.
In the finite element analysis of disc brake squeal GraMQSW16 , the model is a very large-scale second-order system of differential equations with approximately a million degrees of freedom, that furthermore also depends on parameters, e.g., the disc speed $\omega$. If no further constraints are incorporated, then in the stationary case the system takes the form
@@
${\bf J}$ ::=:= $\left[\begin{array}[]{cc}-{\bf G}&-({\bf I}+\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1})\\ ({\bf I}+\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{\bf S}{\bf B}^{\top}){\bf K}^{-1})^{\top}&0\end{array}\right],$
${\bf R}$ ::=:= $\left[\begin{array}[]{cc}{\bf D}&\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1}\\ (\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1})^{\top}&0\end{array}\right],\quad{\bf Q}=\left[\begin{array}[]{cc}{\bf M}^{-1}&0\\ 0&{\bf K}\end{array}\right].$
Since regardless of the choice of ${\bf D}_{\bf S}$, the matrix ${\bf R}$ is indefinite as long as ${\bf N}\neq 0$ (then ${\bf N}-{\bf B}{\bf D}_{\bf S}{\bf B}^{\top}\neq 0$) it is clear that for this system we cannot read off its stability and it is definitely unstable if ${\bf x}^{H}{\bf R}{\bf x}<0$ for some eigenvector ${\bf x}$ of ${\bf J}$.
Remark 29.
Many of the results described in the paper can be extended to the case of general descriptor systems having the form (4) or (7) but with singular ${\bf E}$. This is an open problem and currently a topic of active research.
Remark 30.
In this paper we have always looked for positive definite solutions ${\bf Q}$ of the KYP matrix inequality (9). However, we could relax this requirement to the case that ${\bf Q}$ is positive semidefinite. In this case several further subtle problems arise, and one definitely would need to consider descriptor systems. See CheGH22 ; MehMW18 ; MehS23 for a detailed discussion.
Remark 31.
In view of the fact that the KYP matrix inequality (9) may have many solutions, one may choose the solution so that the resulting pH representation is robust to data or numerical errors. To do this one could choose a robustness measure like to the distance to instability or non-passivity or to consider a representation that is far away from the boundary of the solution set of (9). Partial results in this direction have been obtained in BanMNV20 ; BeaMV19 ; MehV20a .
@@ 5 Numerical Considerations
Another difficulty is the non-uniqueness of solutions to the linear matrix inequality (16). To make the solution unique one can optimize a quality measure like the distance to instability or non-passivity, or try to find the analytic center of the solution, see BanMNV20 ; MehV20 . All these are difficult and expensive optimization problems on top of all the computational work that has to be carried out. How to do this efficiently is an open question, even for the case that the system is minimal and $\hat{\bf S}$ is well-conditioned with respect to inversion. The difficulty arises, in particular, since the solution sets of (16) and (17) are rather difficult to characterize, see MehX24 for a detailed analysis based on eigenvalue perturbation theory.
Remark 32.
Note that if the skew-symmetric/symmetric pencil (106) has purely imaginary eigenvalues, then the solution of the Riccati equation associated with (17) can only be computed with the Newton method of Ben97 which has been implemented e.g. in BenW20 . In this case none of the usual approaches utilizing invariant subspaces will be fully satisfactory and even the Newton method might not achieve a quadratic rate of convergence, displaying only linear convergence GuoL98 .
@@
Remark 33.
[…]
This is common practice when solving (16), see e.g. AndM74 ; HilM80 , where often $\hat{\bf S}$ is perturbed to be invertible, so that the pencil (60) is regular and of index at most one, see Remark 19. However, as discussed in Section 4, see Remark 21, since we are able to deal with a singular $\hat{\bf S}$ we perform the regularization in a different way. For this we have to perform the rank decisions in (62), (77), (97), which are critical in the process. As in most staircase algorithms Van79 , it is recommended to make conservative decisions, i.e., to assume smaller rank if the decision is difficult using the usual rank decision procedures DemK93a ; DemK93b .
Remark 34.
@@
Example 36.
Consider a system (5) with ${\bf A}=({\bf J}-{\bf R}){\bf Q}$, ${\bf B}={\bf F}-{\bf P}$${\bf C}=({\bf F}+{\bf P})^{\top}$${\bf D}={\bf S}+{\bf N}$, where ${\bf J}=0$, ${\bf N}=0$, ${\bf Q}={\bf I}_{4}$,
@@
We ran our script with different values of $\epsilon=0,\pm 2.*e{-51},\ldots,\pm 2.*e{-9}$. When $\epsilon$ is very close to 00 but negative, the system is slightly non-passive, since the matrix ${\bf W}$ has a negative eigenvalue. Our script always returns a positive definite ${\bf Q}$ that changes only slightly as $\epsilon$ varies (regardless of sign), while the inequality (16) remains semidefinite but numerically singular with four very small nonzero eigenvalues. For $\epsilon=0,\pm 10^{-16}$ the pencil (60) has index $\geq 2$, otherwise the index is $1$. Nonetheless, our procedure works in all cases, see Figure 1.
@@
Remark 37.
In Section 4 we have developed a method that allows to identify the solvability of the matrix inequality (9) via the computation of a staircase like form and deflating subspaces of the even matrix pencil (60). In numerical practice, to avoid the use of consecutive rank decisions, one may apply a so called derivative array approach, see KunM24 , where the parts associated with the finite eigenvalues, the part associated with the infinite eigenvalues and the singular part are separated by one sided transformations of (60) from the left generated based on an extended pencil associated with a derivative array of the DAE associated with the even pencil. For even pencils this follows directly from a procedure developed in KunMS13 for time varying DAEs and gives a reduced system that is associated with an even pencil and an unstructured part associated with the eigenvalue $\infty$. We do not present this approach here, see KunMS13 ; KunM23 for details.

### v4

@@ 4.3 Construction in the case of general $\bf D$}
${\bf G}_{1}=\left[\begin{array}[]{rr|r}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}&0\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}&-\Gamma_{1}\\ \hline\cr 0&\Gamma_{1}^{\top}&\Phi_{1}\end{array}\right],$
where $\widetilde{\bf A}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$, $\widetilde{\bf B}_{1},\widetilde{\bf C}_{1}^{\top}\in\mathbb{R}^{n_{1}\times m_{1}}$, $\widetilde{\bf D}_{1}\in\mathbb{R}^{m_{1}\times m_{1}}$, $\Gamma_{1}\in\mathbb{R}^{m_{1}\times d_{1}}$, $\Phi_{1}=-\Phi_{1}^{\top}\in\mathbb{R}^{d_{1}\times d_{1}}$ with $n_{1}=n-r$, $m_{1}=r+s$, and $d_{1}=m-s$. If $\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right]$ satisfies condition (66a), then in an analogous way we construct
[…]
@@
${\bf G}_{2}=\widetilde{\bf T}_{1}\widetilde{\bf V}_{1}^{\top}{\bf G}_{1}\widetilde{\bf V}_{1}\widetilde{\bf T}_{1}^{-1}=\left[\begin{array}[]{rr|rr}-\widetilde{\bf A}_{2}&-\widetilde{\bf B}_{2}&0&0\\ \widetilde{\bf C}_{2}&\widetilde{\bf D}_{2}&-\Gamma_{2}&-\Gamma_{11}\\ \hline\cr 0&\Gamma_{2}^{\top}&\Phi_{2}&-\Gamma_{12}\\ 0&\Gamma_{11}^{\top}&\Gamma_{12}^{\top}&\Phi_{1}\end{array}\right],$
where $\widetilde{\bf A}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$, $\widetilde{\bf B}_{2},\widetilde{\bf C}_{2}^{\top}\in\mathbb{R}^{n_{2}\times m_{2}}$, $\widetilde{\bf D}_{2}\in\mathbb{R}^{m_{2}\times m_{2}}$, $\Gamma_{2}\in\mathbb{R}^{m_{2}\times d_{2}}$, $\Phi_{2}=-\Phi_{2}^{\top}\in\mathbb{R}^{d_{2}\times d_{2}}$ with $n_{2}=n_{1}-r_{1}$, $m_{2}=r_{1}+s_{1}$, $d_{2}=m_{1}-s_{1}$, and $r_{1}=\mathop{\rm rank}\nolimits\widetilde{\bf B}_{1}$, $s_{1}=\mathop{\rm rank}\nolimits(\widetilde{\bf D}_{1}+\widetilde{\bf D}_{1}^{\top})$. We then proceed.
The procedure stops in two cases, (a) condition (66a) is violated or $\widetilde{\bf D}_{j}+\widetilde{\bf D}_{j}^{\top}$ is indefinite for some $j$ and no transformation to port-Hamiltonian form exists, or (b) after $k$ steps one arrives at
@@
[…]
and $n_{j}$$m_{j}$ are the sizes of ${\bf T}_{j}$, ${\bf V}_{j}$, respectively, for $j=0,\ldots,k-1$. Case (b) happens because $n_{j}$ is strictly decreasing until $r_{j}$ is void or zero for some $j$. If $r_{j}$ is void, then $\widetilde{\bf D}_{j}+\widetilde{\bf D}_{j}$ is positive definite already. If $r_{j}=0$, then $n_{j+1}=n_{j}$ and $m_{j+1}=s_{j}$ indicating that $\widetilde{\bf D}_{j+1}+\widetilde{\bf D}_{j+1}^{\top}$ is positive definite.
@@
Remark 27
The recursive procedure described above requires at each step the computation of three singular value decompositions in order to check the ranks of the matrices $\widetilde{\bf B}_{j}$ and $\widetilde{\bf C}_{j}$ and in order to construct bi-orthogonal bases so that (80) holds. While each step of this procedure can be implemented in a numerically stable way, the consecutive rank decisions make the aggregate procedure difficult to analyze, similar to the case of staircase algorithms [15, 20, 21]. In general the strategy should be adapted toward the goal of obtaining a realization in port-Hamiltonian form that is robust to small perturbations, see [4, 46] for some ways to do this.
@@ 4.4 Explicit solution of linear matrix inequalities via even pencils
We have seen that to check the existence of the transformation to port-Hamiltonian form and to explicitly construct the transformation matrices ${\bf T},{\bf V}$ in (8) is equivalent to considering the solution of the linear matrix inequality (9). In this subsection we combine the recursive procedure presented in the previous subsection with the construction of a staircase like form ([62]) for the even pencil (60).
@@
[…]
$=:$ $\left[\begin{array}[]{cc}-{\bf A}_{1}^{\top}{\bf Q}_{1}-{\bf Q}_{1}{\bf A}_{1}&{\bf C}_{1}^{\top}-{\bf Q}_{1}{\bf B}_{1}\\ {\bf C}_{1}-{\bf B}_{1}^{\top}{\bf Q}_{1}&{\bf D}_{1}+{\bf D}_{1}^{\top}\end{array}\right]\geq 0$
@@
Note the above process is actually a special staircase form reduction process that deflates the singular part and higher index of the eigenvalue infinity of the even pencil (60), [13, 15].
@@
If we consider the positive definite solution of the KYP equation, we can directly check if (106) has a deflating subspace associated with a set of $\ell$ finite eigenvalues chosen such that the deflating subspace is as in (60). If such a deflating subspace exists and ${\bf W}_{1}$ is invertible, then we can compute a Hermitian positive definite matrix ${\bf Q}_{k}$ associated with (106). We comment that an analogous characterization of solutions of singular H control problems via matrix pencils was given in [19].
Remark 28
Note that to check the passivity of (5) it is only necessary to have a positive semidefinite solution to (9). Thus, if one only wants to check passivity, then Part a) in Lemma 23 can be relaxed to $\mathsf{Ker}\,{\bf B}_{1}\subseteq\mathsf{Ker}\,{\bf C}_{1}^{\top}$. In this case the transformation to the form (97) can still be made, but ${\bf Q}_{22}$ in (98) is only positive semidefinite, and $\mathsf{Ker}\,{\bf Q}_{22}\subseteq\mathsf{Ker}\,{\bf Q}_{12}$. Then ${\bf Q}_{0}$ can still be defined but instead of ${\bf Q}_{11}^{-1}$ one needs to use the Moore-Penrose pseudoinverse, see [25], of ${\bf Q}_{11}$. However, in this case ${\bf Q}_{0}$ and the resulting solution $\bf Q$} cannot be positive definite. Thus, in this situation, (5) may be a passive system that cannot be transformed to a standard port-Hamiltonian system of the form (6) system.
@@
Example 29
In the finite element analysis of disc brake squeal [26], the model is a very large-scale second-order system of differential equations with approximately a million degrees of freedom, that furthermore also depends on parameters, e.g., the disc speed $\omega$. If no further constraints are incorporated, then in the stationary case the system takes the form
@@
${\bf J}$ $:=$ $\left[\begin{array}[]{cc}-{\bf G}&-({\bf I}+\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1})\\ ({\bf I}+\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{\bf S}{\bf B}^{\top}){\bf K}^{-1})^{\top}&0\end{array}\right],$
${\bf R}$ $:=$ $\left[\begin{array}[]{cc}{\bf D}&\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1}\\ (\frac{1}{2}({\bf N}-{\bf B}{\bf D}_{{\bf S}}{\bf B}^{\top}){\bf K}^{-1})^{\top}&0\end{array}\right],\quad{\bf Q}=\left[\begin{array}[]{cc}{\bf M}^{-1}&0\\ 0&{\bf K}\end{array}\right].$
Regardless of the choice of the matrix ${\bf D}_{\bf S}$ in the feedback $\bf u$}, the matrix $\bf R$} is indefinite as long as ${\bf N}\neq 0$ (then ${\bf N}-{\bf B}{\bf D}_{\bf S}{\bf B}^{\top}\neq 0$); thus, it is clear that for this system we cannot read off its stability and it is definitely unstable if ${\bf x}^{H}{\bf R}{\bf x}<0$ for some eigenvector $\bf x$} of $\bf J$}. This shows that the rewriting of $\bf A$} as $({\bf J}-{\bf R}){\bf Q})$ alone will not be enough to check stability, but if the system is asymptotically stable then a further transformation as described here will be necessary.
Remark 30
Many of the results described in this work can be extended to the case of general descriptor systems having the form (11) or (7) but with singular $\bf E$}. This is an open problem and currently a topic of active research.
Remark 31
In this paper, we only consider positive definite solutions $\bf Q$} of the KYP matrix inequality (9). In principle, this condition could be relaxed to requiring that $\bf Q$} is only positive semidefinite. However, several further subtle problems arise in this case, and many of them lead to the need to consider descriptor systems. See [18, 43, 49] for a detailed discussion.
Remark 32
In view of the fact that the KYP matrix inequality (9) may have many solutions, one may choose the solution so that the resulting pH representation is robust to data or numerical errors. To do this one could choose a robustness measure like to the distance to instability or non-passivity or to consider a representation that is far away from the boundary of the solution set of (9). Partial results in this direction have been obtained in [4, 6, 45].
@@ 5 Numerical Considerations
Another difficulty is the non-uniqueness of solutions to the linear matrix inequality (16). To make the solution unique one can optimize a quality measure like the distance to instability or non-passivity, or try to find the analytic center of the solution, see [4, 46]. All these are difficult and expensive optimization problems on top of all the computational work that has to be carried out. How to do this efficiently is an open question, even for the case that the system is minimal and $\hat{\bf S}$ is well-conditioned with respect to inversion. The difficulty arises, in particular, since the solution sets of (16) and (17) are rather difficult to characterize, see [51] for a detailed analysis based on eigenvalue perturbation theory.
Remark 33
Note that if the skew-symmetric/symmetric pencil (106) has purely imaginary eigenvalues, then the solution of the Riccati equation associated with (17) can only be computed with the Newton method of [8] which has been implemented e.g. in [11]. In this case none of the usual approaches utilizing invariant subspaces will be fully satisfactory and even the Newton method might not achieve a quadratic rate of convergence, displaying only linear convergence [29].
@@
Remark 34
As we have discussed in Section 4, we reduce the problem to a subproblem, where $\hat{\bf C}$ is invertible and thus where (17) can be formed and is solvable. For this we have to make several rank decisions or regularization steps which result in small perturbations. This is often justified, since the coefficient matrices $\bf A$}, $\bf B$}, $\bf C$}, $\bf D$} are typically not exact because they arise often from a data based realization, interpolation or model reduction process. So we can make small perturbations to these data to regularize the problem if this does not change the resulting (hopefully) robust pH representation.
This is common practice when solving (16), see e.g. [3, 30], where often $\hat{\bf S}$ is perturbed to be invertible, so that the pencil (60) is regular and of index at most one, see Remark 19. However, as discussed in Section 4, see Remark 21, since we are able to deal with a singular $\hat{\bf S}$ we perform the regularization in a different way. For this we have to perform the rank decisions in (62), (77), (97), which are critical in the process. As in most staircase algorithms [62], it is recommended to make conservative decisions, i.e., to assume smaller rank if the decision is difficult using the usual rank decision procedures [20, 21].
Remark 35
@@
Example 37
Consider a system (5) with ${\bf A}=({\bf J}-{\bf R}){\bf Q}$, ${\bf B}={\bf F}-{\bf P}$, ${\bf C}=({\bf F}+{\bf P})^{\top}$, ${\bf D}={\bf S}+{\bf N}$, where ${\bf J}=0$, ${\bf N}=0$, ${\bf Q}={\bf I}_{4}$,
@@
We ran our script with different values of $\epsilon=0,\pm 2.*e{-51},\ldots,\pm 2.*e{-9}$. When $\epsilon$ is very close to 0 but negative, the system is slightly non-passive, since the matrix $\bf W$} has a negative eigenvalue. Our script always returns a positive definite $\bf Q$} that changes only slightly as $\epsilon$ varies (regardless of sign), while the inequality (16) remains semidefinite but numerically singular with four very small nonzero eigenvalues. For $\epsilon=0,\pm 10^{-16}$ the pencil (60) has index $\geq 2$, otherwise the index is $1$. Nonetheless, our procedure works in all cases, see Figure 1, where by ’residual’ we denote the norm of the perturbation that we have to add to the linear matrix inequality (9) to guarantee the existence of a port-Hamiltonian formulation.
@@
Remark 38
In Section 4 we have developed a method that allows to identify the solvability of the matrix inequality (9) via the computation of a staircase like form and deflating subspaces of the even matrix pencil (60). In numerical practice, to avoid the use of consecutive rank decisions, one may apply a so called derivative array approach, see [38], where the parts associated with the finite eigenvalues, the part associated with the infinite eigenvalues and the singular part are separated by one sided transformations of (60) from the left generated based on an extended pencil associated with a derivative array of the DAE associated with the even pencil. For even pencils this follows directly from a procedure developed in [39] for time varying DAEs and gives a reduced system that is associated with an even pencil and an unstructured part associated with the eigenvalue $\infty$. We do not present this approach here, see [39, 37] for details.